# Finite-Precision Conjugate Gradient Method for Massive MIMO Detection

### v1

@@ Abstract
Index Terms:
Block Jacob, conjugate gradient, finite-precision arithmetic, massive MIMO detection
@@ I Introduction
In addition to the hardware complexity, the computational complexity of massive MIMO detection has also generated considerable interest. Even for the linear detection schemes like ZF and LMMSE, the increasing system dimension leads to the complicated matrix inversion. To address this problem, various iterative MIMO detection methods have been proposed, which can be broadly classified into matrix-splitting (MS)/stationary methods and gradient-based/non-stationary methods. In MS methods, the Gram matrix is typically decomposed into its diagonal, upper, and lower triangular components. Specifically, the Jacobi iteration (JI) method [8] and the Gauss-Seidel (GS) method [9] used the inversion of the diagonal and lower triangular parts, respectively. As for the successive over-relaxation (SOR) method [10], the inversion of both the upper and lower triangular parts are incorporated to achieve faster convergence than the JI and GS methods. Moreover, the authors in [11] proposed the randomized iterative detection algorithm that adopts the random sampling into the MS methods.
Moreover, gradient-based iterative methods are promising to reduce the computational complexity of massive MIMO detection. The authors in [12] employed gradient descent (GD) to implement iterative MIMO detection. Besides, the steepest descent (SD) method [13] and Barzilai-Bowein (BB) method [14] were introduced to solve the problem of massive MIMO detection. The authors in [15] leveraged limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) to massive MIMO detection, demonstrating convergence comparable to that of BFGS but with lower complexity. In addition, the conjugate gradient (CG) method exploits the Hermitian property of the Gram matrix to determine more efficient update directions [16], thereby accelerating convergence compared to GD, SD, and BB methods while achieving performance equivalent to L-BFGS.
@@
The aforementioned works on low-complexity CG detection commonly implement the matrix computations using full-precision arithmetic ($\mathtt{fp64}$). Note that finite-precision or low-precision arithmetic offers further opportunities to reduce computational demands compared with full-precision arithmetic [24]. For instance, half-precision arithmetic provides advantages over double-precision arithmetic, including a fourfold increase in processing speed, a quarter of the storage requirements, and a $75\%$ reduction in memory transfer costs [25]. Recently, finite-precision arithmetic has emerged as a powerful technique for reducing computational complexity in computer science and wireless communications. Specifically, large language models (LLMs), such as Deepseek [26, 27] and LLaMA [28, 29], frequently utilize finite-precision arithmetic to accelerate matrix computations. In the context of wireless communications, the impact of finite-precision arithmetic on communication performance was analyzed, and a mixed-precision transceiver design was proposed in [30].
@@ II System Model
where $\zeta$ is the correlation coefficient with $10\leq\zeta<$ 1.
Then, to recover the transmitted signals, classic linear detection like ZF and LMMSE is employed, which can be written as [32]
@@ 9      Update ${\bf p}_{i}$ using (10)
Consider a multi-user MIMO (MU-MIMO) system using full-precision arithmetic ($\mathtt{fp64}$), where the BS is equipped with a large number of antennas, e.g., $M=512$, serving substantial single-antenna users, e.g., $N=K=100$. According to the Monte-Carlo simulation, CG detection requires approximately $28$ iterations to achieve the BER performance of LMMSE detection at a high SNR level (SNR $=25$ dB) and under highly correlated channels ($\zeta=0.8$). This corresponds to an estimated computational cost of approximately 280,000 floating-point operations (flops) per subchannel. Further, considering the orthogonal frequency-division multiplexing (OFDM) system supporting up to $3300$ subcarriers with a slot duration of $30\mathrm{\mu s}$ [34], the required computational cost for CG detection is on the order of $30$ tera floating-point operations per second (TFLOPS). Notably, the NVIDIA A100 graphics processing unit (GPU) can deliver over $9.7$ TFLOPS [35]. Consequently, multiple GPUs would be required to meet the computational demand, leading to an impractical increase in computational cost and power consumption.
@@ III Finite-Precision CG Detection
In this section, we propose an FP-CG detection to mitigate computational bottleneck in Algorithm 1. First, we introduce the detailed FP-CG detection. Then, the performance of FP-CG detection is analyzed to evaluate the impact of finite-precision arithmetic, and we also give a heuristic method to show how to choose different precisions in different channel conditions without significantly affecting performance.
@@ III-A Proposed FP-CG Detection
$\left(\mathrm{sig.},\mathrm{exp.}\right)$(1) $u$(2) $x_{\min}$(3) $x_{\max}$(4)
@@
Denote by $u$ the unit roundoff. The floating-point system $\mathbb{F}$ adheres to a standard arithmetic model if, for any $x,y\in\mathbb{R}$ is in the range of $\mathbb{F}$. One has[24, Sec. 2.2]
@@
The standard arithmetic model in Definition 2 is satisfied by the IEEE arithmetic standard [25]. Various unit roundoffs $u$ in Table I can be chosen to represent different precisions.
@@ III-B Performance Analysis of FP-CG Detection
Lemma 1 (Real-valued Matrix Computations [24, 36]).
@@
The proof is similar to that of [30, Theorem 1 & 2], which is omitted for conciseness. ∎
Then, based on Theorem 1, we can analyze the maximum attainable accuracy of FP-CG detection in Algorithm 2. Due to finite-precision arithmetic, the true residuals ${\bf b-Ax}_{i}$ and the computed residuals vectors ${\bf r}_{i}$ in Algorithm 2 have a gap. Following the work of [36], we can use the difference between ${\bf b-Ax}_{i}$ and ${\bf r}_{i}$ to analyze the maximum attainable accuracy of Algorithm 2, which is derived in the following theorem.
@@
Theorem 2 reveals that the attainable accuracy of Algorithm 2 depends on work precision $u_{\rm W}$, matrix-vector products precision $u_{\rm MV}$, the dimension of ${\bf A}$, and condition number $\kappa\left(\mathbf{A}\right)$. Under real-valued computations with $u_{\rm W}=u_{\rm MV}=u_{\rm IP}$, (14) reduces to the special case of real-valued CG using uniform precision shown in [36, Eq. (24)]. The insights from Theorem 2 are further highlighted in the following remark.
@@
In Theorem 2, (14) is derived under the assumption of convergence, and hence we need to discuss the convergence properties of Algorithm 2. On the one hand, using finite-precision arithmetic, particularly low-precision arithmetic, may delay convergence due to the loss of local orthogonality [37]. On the other hand, applying the conclusion from [38, Theorem 5.24] to Algorithm 2 reveals that if $u_{\mathrm{MV}}\left\|\mathbf{A}^{-1}\right\|_{2}\ll 1$, convergence is guaranteed.
As discussed in Remark 1 and 2, employing finite-precision arithmetic, especially low-precision arithmetic, can negatively impact the attainable accuracy and convergence of CG detection. This raises a critical question: How should one select precisions to avoid significant degradation in CG detection performance? To this end, we expect that
@@
Since the derivation of (18) is based on Theorem 1, which provides worst-case bounds, these bounds tend to be conservative and overestimated [39]. Therefore, we can roughly omit $\mathcal{O}(\cdot)$, replace $\ll$ with $<$, and formulate the heuristic method:
@@
The computational complexity of FP-CG detection is significantly lower than that of CG detection due to the utilization of finite-precision arithmetic. For instance, in non-correlated ($\zeta=0$) and moderately correlated channels ($\zeta=0.5$), FP-CG detection can reduce computational cost by $75\%$ compared with CG detection (See Fig. 2a and 2b). In high correlated channels ($\zeta=0.8$), FP-CG detection achieves a $43.3\%$ reduction in computational cost compared to CG detection (See Fig. 2c and 6a).
@@ IV Joint Finite Precision and Preconditioning CG Detection
In this section, we consider decreasing the iteration number based on FP-CG detection to further reduce the computational complexity of CG detection. First, the BJ preconditioning is introduced based on the correlated channel property, and then FP-BJ-CG detection is proposed. Finally, we present the performance analysis of FP-BJ-CG detection.
@@ IV-A Block Jacobi Preconditioning and FP-BJ-CG Detection
Theorem 3 shows that the Gram matrix of correlated channel asymptotically approaches deterministic exponential correlation matrices $\mathbf{R}_{t}$. Since inter-user correlation is typically weak [20], the Gram matrix tends to be near block diagonally dominant for massive MIMO systems [^2] This behavior has been observed in [43, 44] by simulation. To our knowledge, this paper provides the first theoretical proof of this behavior., and the block property becomes more evident with the increase of correlation coefficient $\zeta$. Furthermore, when $\zeta=0$, Theorem 3 can reduce to the well-known favorable propagation of Rayleigh fading channel in [45].
Based on the near block diagonally dominant property, the BJ preconditioning [46] is introduced for joint finite precision and precondition CG detection. This can further reduce the computational complexity of CG detection by lowering the iteration count required for convergence. Specifically, the preconditioning matrix ${\bf P}_{d}$ uses the block diagonal entries of ${\bf A}$ and contains $d$ block sub-matrices. For simplification, the dimensions of block sub-matrices are assumed to be identical, denoted as $L=N/d\in\mathbb{Z}$ with $L<N$. Then the specific expression of ${\bf P}_{d}\in\mathbb{C}^{N\times N}$ and its inversion $\mathbf{M}_{d}\in\mathbb{C}^{N\times N}$ are given by
@@ 34         $\triangleright$ $u_{\rm W}$
Given the BJ preconditioning, we present a basic implementation of FP-BJ-CG detection in Algorithm 3. More precisely, Algorithm 3 employs a left-preconditioned CG method, which implicitly uses the asymmetric matrix ${\bf M}_{d}{\bf A}$ instead of the symmetric matrix $\bar{\bf A}={\bf M}_{d}^{1/2}{\bf A}{\bf M}_{d}^{1/2}$. Both approaches produce identical sequences of iterates [47], yet the former is more computationally efficient. For convenience, the following theoretical analysis results are presented regarding the symmetrically preconditioned matrix, i.e., $\bar{\bf A}$.
@@ IV-B Performance Analysis of FP-BJ-CG Detection
Theorem 4 demonstrates that the attainable accuracy of FP-BJ-CG detection depends on work precision $u_{\rm W}$, matrix-vector products precision $u_{\rm MV}$, the dimension of $\bar{\bf A}$, and condition number $\kappa\left(\bar{\bf A}\right)$. Furthermore, similar to the analysis of Theorem 2, we can obtain the same results of Remark 1 and 2. Moreover, the heuristic method is changed into
@@
$\lim_{M\rightarrow\infty}{\bf H}_{k}^{H}{\bf H}_{j}=0,~{}\forall k\neq j.$ (31)
@@
The proof is available in Appendix C. ∎
@@
Finally, we present the computational complexity of FP-BJ-CG detection. The computational complexity of FP-BJ-CG detection comprises two components: preconditioning and the iterative process. For the preconditioning process, the computational complexity is mainly caused by the inversion of block diagonal matrix ${\bf P}_{d}$, leading to $d\times\mathcal{O}\left(L^{3}\right)=\mathcal{O}\left(NL^{2}\right)$. In the iterative process, since ${\bf M}_{d}$ is block diagonal, the computational complexity of ${\bf M}_{d}{\bf r}_{i}$ in Step 10 is $\mathcal{O}\left(NL\right)$, still lower than that of ${\bf A}{\bf p}_{i-1}$ in Step 4. Therefore, ${\bf A}{\bf p}_{i-1}$ dominant the computational complexity of the iterative process and requires $\mathcal{O}\left(N^{2}\right)$. In summary, the overall computational complexity of FP-BJ-CG detection in Algorithm 3 is $\mathcal{O}\left(NL^{2}+\mathcal{I}N^{2}\right)$.
Remark 3 (Computational Complexity Reduction).
@@
Despite the extra computational complexity of BJ preconditioning, FP-BJ-CG detection can reduce the computational complexity by decreasing the iteration number and supporting lower-precision arithmetic. For example, considering $N_{\rm UE}=4$, $N=32$, and $L=8$, the extra computational complexity of BJ preconditioning is equivalent to $2$ CG detection iteration. Notably, Fig. 3 shows that FP-BJ-CG detection can accelerate CG detection by up to approximately $5$ iterations and support half-precision arithmetic ($\mathtt{fp16}$) to achieve the accuracy of LMMSE detection. Specifically, FP-BJ-CG detection with $\mathtt{fp16}$ decreases $40.0\%$ and $81.3\%$ computational cost compared to CG and LMMSE detection with $\mathtt{fp64}$, respectively (See Fig. 6b). These advantages of FP-BJ-CG detection significantly offset its additional preconditioning cost and reduce the overall computational complexity.

### v2

@@ I Introduction
In addition to the hardware complexity, the computational complexity of massive MIMO detection has also generated considerable interest. Even for the linear detection schemes like ZF and LMMSE, the increasing system dimension leads to complicated matrix inversion. To address this problem, various iterative MIMO detection methods have been proposed, which can be broadly classified into matrix-splitting (MS)/stationary methods and gradient-based/non-stationary methods. In MS methods, the Gram matrix is typically decomposed into its diagonal, upper, and lower triangular components. Specifically, the Jacobi iteration (JI) method [7794623] and the Gauss-Seidel (GS) method [8511052] used the inversion of the diagonal and lower triangular parts, respectively. As for the successive over-relaxation (SOR) method [7399337], the inversion of both the upper and lower triangular parts are incorporated to achieve faster convergence than the JI and GS methods. Moreover, the authors in [9790338] proposed the randomized iterative detection algorithm that adopts random sampling into the MS methods.
Moreover, gradient-based iterative methods are promising to reduce the computational complexity of massive MIMO detection. The authors in [10024792] employed gradient descent (GD) to implement iterative MIMO detection. Besides, the steepest descent (SD) method [xue2018adaptive] and Barzilai-Bowein (BB) method [8598346] were introduced to solve the problem of massive MIMO detection. The authors in [9806305] leveraged limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) to massive MIMO detection, demonstrating convergence comparable to that of BFGS but with lower complexity. In addition, the conjugate gradient (CG) method exploits the Hermitian property of the Gram matrix to determine more efficient update directions [7037382], thereby accelerating convergence compared to GD, SD, and BB methods while achieving performance equivalent to L-BFGS.
@@
The aforementioned works on low-complexity CG detection commonly implement the matrix computations using full-precision arithmetic ($\mathtt{fp64}$). Note that finite-precision or low-precision arithmetic offers further opportunities to reduce computational demands compared with full-precision arithmetic [higham2002accuracy]. For instance, half-precision arithmetic provides advantages over double-precision arithmetic, including a fourfold increase in processing speed, a quarter of the storage requirements, and a $75\%$ reduction in memory transfer costs [higham2022mixed]. Recently, finite-precision arithmetic has emerged as a powerful technique for reducing computational complexity in computer science and wireless communications. Specifically, large language models (LLMs), such as Deepseek [liu2024deepseek, liu2024deepseek1] and LLaMA [touvron2023llama, 10884770], frequently utilize finite-precision arithmetic to accelerate matrix computations. In the context of wireless communications, the authors in [7168929] and [9110827] proposed fixed-point arithmetic CG detection and finite-alphabet MIMO detection, respectively. Moreover, the impact of finite-precision arithmetic on communication performance was analyzed, and a mixed-precision transceiver design was proposed in [10845867].
@@ II System Model
where $\zeta$ is the correlation coefficient with $0\leq\zeta<1$.
Then, to recover the transmitted signals, classic linear detection like ZF and LMMSE is employed, which can be written as [yang2015fifty]
@@ 9  Update ${\bf p}_{i}$ using (10)
Consider a multi-user MIMO (MU-MIMO) system using full-precision arithmetic ($\mathtt{fp64}$), where the BS is equipped with a large number of antennas, e.g., $M=512$, serving substantial single-antenna users, e.g., $N=K=100$. According to the Monte-Carlo simulation, CG detection requires approximately $28$ iterations to achieve the BER performance of LMMSE detection at a high SNR level (SNR =25=25 dB) and under highly correlated channels ($\zeta=0.8$). This corresponds to an estimated computational cost of approximately 280,000 floating-point operations (flops) per subchannel. Further, considering the orthogonal frequency-division multiplexing (OFDM) system supporting up to $3300$ subcarriers with a slot duration of $30\mathrm{\mu s}$ [8412469], the required computational cost for CG detection is on the order of $30$ tera floating-point operations per second (TFLOPS). Notably, the NVIDIA A100 graphics processing unit (GPU) can deliver over $9.7$ TFLOPS [nvidia]. Consequently, multiple GPUs would be required to meet the computational demand, leading to an impractical increase in computational cost and power consumption.
@@ III Finite-Precision CG Detection
In this section, we propose an FP-CG detection to mitigate computational bottlenecks in Algorithm 1. First, we introduce the detailed FP-CG detection. Then, the performance of FP-CG detection is analyzed to evaluate the impact of finite-precision arithmetic, and we also give a heuristic method to show how to choose different precisions in different channel conditions without significantly affecting performance.
@@ III-A Proposed FP-CG Detection
$\left(\mathrm{sig.},\mathrm{exp.}\right)$(1) uu(2) $x_{\min}$(3) $x_{\max}$(4)
@@
Denote by $u$ the unit roundoff. The floating-point system $\mathbb{F}$ adheres to a standard arithmetic model if, for any $x,y\in\mathbb{R}$ is in the range of $\mathbb{F}$. One has[higham2002accuracy, Sec. 2.2]
@@
The standard arithmetic model in Definition 2 is satisfied by the IEEE arithmetic standard [higham2022mixed]. Various unit roundoffs $u$ in Table I can be chosen to represent different precisions.
@@ III-B Performance Analysis of FP-CG Detection
Lemma 1 (Real-valued Matrix Computations [higham2002accuracy, greenbaum1997estimating]).
@@
The proof is similar to that of [10845867, Theorem 1 & 2], which is omitted for conciseness. ∎
Then, based on Theorem 1, we can analyze the maximum attainable accuracy of FP-CG detection in Algorithm 2. Due to finite-precision arithmetic, the true residuals ${\bf b-Ax}_{i}$ and the computed residuals vectors ${\bf r}_{i}$ in Algorithm 2 have a gap. Following the work of [greenbaum1997estimating], we can use the difference between ${\bf b-Ax}_{i}$ and ${\bf r}_{i}$ to analyze the maximum attainable accuracy of Algorithm 2, which is derived in the following theorem.
@@
Theorem 2 reveals that the attainable accuracy of Algorithm 2 depends on work precision $u_{\rm W}$, matrix-vector products precision $u_{\rm MV}$, the dimension of $\bf A$}, and condition number $\kappa\left(\mathbf{A}\right)$. Under real-valued computations with $u_{\rm W}=u_{\rm MV}=u_{\rm IP}$, (14) reduces to the special case of real-valued CG using uniform precision shown in [greenbaum1997estimating, Eq. (24)]. The insights from Theorem 2 are further highlighted in the following remark.
@@
In Theorem 2, (14) is derived under the assumption of convergence, and hence we need to discuss the convergence properties of Algorithm 2. On the one hand, using finite-precision arithmetic, particularly low-precision arithmetic, may delay convergence due to the loss of local orthogonality [greenbaum2021convergence]. On the other hand, applying the conclusion from [meurant2006lanczos, Theorem 5.24] to Algorithm 2 reveals that if $u_{\mathrm{MV}}\left\|\mathbf{A}^{-1}\right\|_{2}\ll 1$, convergence is guaranteed.
As discussed in Remarks 1 and 2, employing finite-precision arithmetic, especially low-precision arithmetic, can negatively impact the attainable accuracy and convergence of CG detection. This raises a critical question: How should one select precisions to avoid significant degradation in CG detection performance? To this end, we expect that
@@
Since the derivation of (18) is based on Theorem 1, which provides worst-case bounds, these bounds tend to be conservative and overestimated [higham2019new]. Therefore, we can roughly omit $\mathcal{O}(\cdot)$, replace $\ll$ with <<, and formulate the heuristic method:
@@
The computational complexity of FP-CG detection is significantly lower than that of CG detection due to the utilization of finite-precision arithmetic. For instance, in non-correlated ($\zeta=0$) and moderately correlated channels ($\zeta=0.5$), FP-CG detection can reduce computational cost by $75\%$ compared with CG detection (See Figs. 2a and 2b). In highly correlated channels ($\zeta=0.8$), FP-CG detection achieves a $43.3\%$ reduction in computational cost compared to CG detection (See Figs. 2c and 7a).
@@ IV Joint Finite Precision and Preconditioning CG Detection
In this section, we consider decreasing the iteration number based on FP-CG detection to further reduce the computational complexity of CG detection. First, the BJ preconditioning is introduced based on the correlated channel property, and then the FP-BJ-CG detection is proposed. Finally, we present the performance analysis of FP-BJ-CG detection.
@@ IV-A Block Jacobi Preconditioning and FP-BJ-CG Detection
Theorem 3 shows that the Gram matrix of correlated channel asymptotically approaches deterministic exponential correlation matrices $\mathbf{R}_{t}$. Since inter-user correlation is typically weak [10653741], the Gram matrix tends to be near block diagonally dominant for massive MIMO systems [^2] This behavior has been observed in [bentrcia2019domain, 9374779] by simulation. To our knowledge, this paper provides the first theoretical proof of this behavior., and the block property becomes more evident with the increase of correlation coefficient $\zeta$. Furthermore, when $\zeta=0$, Theorem 3 can reduce to the well-known favorable propagation of the Rayleigh fading channel in [6951994].
Based on the near block diagonally dominant property, the BJ preconditioning [eisenstat1981efficient] is introduced for joint finite precision and precondition CG detection (Appendix C compares the BJ preconditioning with other preconditioning methods). This can further reduce the computational complexity of CG detection by lowering the iteration count required for convergence. Specifically, the preconditioning matrix ${\bf P}_{d}$ uses the block diagonal entries of $\bf A$} and contains $d$ block sub-matrices. For simplification, the dimensions of block sub-matrices are assumed to be identical, denoted as $L=N/d\in\mathbb{Z}$ with $L<N$. Then the specific expression of ${\bf P}_{d}\in\mathbb{C}^{N\times N}$ and its inversion $\mathbf{M}_{d}\in\mathbb{C}^{N\times N}$ are given by
@@ 34  $\triangleright$ $u_{\rm W}$
Given the BJ preconditioning, we present a basic implementation of FP-BJ-CG detection in Algorithm 3. More precisely, Algorithm 3 employs a left-preconditioned CG method, which implicitly uses the asymmetric matrix ${\bf M}_{d}{\bf A}$ instead of the symmetric matrix $\bar{\bf A}={\bf M}_{d}^{1/2}{\bf A}{\bf M}_{d}^{1/2}$. Both approaches produce identical sequences of iterates [saad2003iterative], yet the former is more computationally efficient. For convenience, the following theoretical analysis results are presented regarding the symmetrically preconditioned matrix, i.e., $\bar{\bf A}$.
@@ IV-B Performance Analysis of FP-BJ-CG Detection
Theorem 4 demonstrates that the attainable accuracy of FP-BJ-CG detection depends on work precision $u_{\rm W}$, matrix-vector products precision $u_{\rm MV}$, the dimension of $\bar{\bf A}$, and condition number $\kappa\left(\bar{\bf A}\right)$. Furthermore, similar to the analysis of Theorem 2, we can obtain the same results of Remarks 1 and 2. Moreover, the heuristic method is changed into
@@
$\lim_{M\rightarrow\infty}{\bf H}_{k}^{\mathsf{H}}{\bf H}_{j}=0,~\forall k\neq j.$ (31)
@@
The proof is available in Appendix D. ∎
@@
Finally, we present the computational complexity of FP-BJ-CG detection. The computational complexity of FP-BJ-CG detection comprises two components: preconditioning and the iterative process. For the preconditioning process, the computational complexity is mainly caused by the inversion of the block diagonal matrix ${\bf P}_{d}$, leading to $d\times\mathcal{O}\left(L^{3}\right)=\mathcal{O}\left(NL^{2}\right)$. In the iterative process, since ${\bf M}_{d}$ is block diagonal, the computational complexity of ${\bf M}_{d}{\bf r}_{i}$ in Step 10 is $\mathcal{O}\left(NL\right)$, still lower than that of ${\bf A}{\bf p}_{i-1}$ in Step 4. Therefore, ${\bf A}{\bf p}_{i-1}$ dominant the computational complexity of the iterative process and requires $\mathcal{O}\left(N^{2}\right)$. In summary, the overall computational complexity of FP-BJ-CG detection in Algorithm 3 is $\mathcal{O}\left(NL^{2}+\mathcal{I}N^{2}\right)$.
Remark 3 (Computational Complexity Reduction Compared with CG Detection).
@@
Despite the extra computational complexity of BJ preconditioning, FP-BJ-CG detection can reduce the computational complexity by decreasing the iteration number and supporting lower-precision arithmetic. For example, considering $N_{\rm UE}=4$, $N=32$, and $L=8$, the extra computational complexity of BJ preconditioning is equivalent to $2$ CG detection iterations. Notably, Fig. 3 shows that FP-BJ-CG detection can accelerate CG detection by up to approximately $5$ iterations and support half-precision arithmetic ($\mathtt{fp16}$) to achieve the accuracy of LMMSE detection. Specifically, FP-BJ-CG detection with $\mathtt{fp16}$ decreases $30.0\%$ and $78.2\%$ computational cost compared to CG and LMMSE detection with $\mathtt{fp64}$, respectively (See Fig. 7b). These advantages of FP-BJ-CG detection significantly offset its additional preconditioning cost and reduce the overall computational complexity.
Remark 4 (Trade-off Analysis between FP-CG and FP-BJ-CG Detection).
The computational complexity of the preconditioning step is $\mathcal{O}\left(NL^{2}\right)$. Although this represents a one-time cost per channel realization, it may constitute a non-trivial overhead compared with FP-CG detection. To quantitatively assess this trade-off, we derive a general break-even criterion that identifies when preconditioning becomes computationally advantageous. Let $C_{\mathrm{BJ}}$ denote the preconditioning setup cost and let $C_{\mathrm{PCG}}^{\mathrm{it}}$ and $C_{\mathrm{CG}}^{\mathrm{it}}$ denote the per-iteration computational cost of the preconditioned and non-preconditioned FP-CG methods, respectively. Assuming FP-BJ-CG and FP-CG use the same finite-precision arithmetic, the total costs are
$C_{\mathrm{PCG}}=C_{\mathrm{BJ}}+k_{\mathrm{PCG}}C_{\mathrm{PCG}}^{\mathrm{it}},~C_{\mathrm{CG}}=k_{\mathrm{CG}}C_{\mathrm{CG}}^{\mathrm{it}},$ (34)
where $k_{\mathrm{PCG}}$ and $k_{\mathrm{CG}}$ are the number of iterations by the preconditioned and non-preconditioned algorithms to reach the target accuracy. Notably, the preconditioning becomes computationally advantageous when $C_{\mathrm{PCG}}<C_{\mathrm{CG}}$, i.e.,
$k_{\mathrm{CG}}-k_{\mathrm{PCG}}>\frac{C_{\mathrm{BJ}}}{C_{\mathrm{CG}}^{\mathrm{it}}}+k_{\mathrm{PCG}}\left(\frac{C_{\mathrm{PCG}}^{\mathrm{it}}}{C_{\mathrm{CG}}^{\mathrm{it}}}-1\right).$ (35)
Since the dominant per-iteration cost is the same matrix–vector product for both methods, we have $C_{\mathrm{PCG}}^{\mathrm{it}}\approx C_{\mathrm{CG}}^{\mathrm{it}}\triangleq C^{\mathrm{it}}$. Under this assumption, the break-even condition simplifies to
$\Delta k\triangleq k_{\mathrm{CG}}-k_{\mathrm{PCG}}>\frac{C_{\mathrm{BJ}}}{C^{\mathrm{it}}}.$ (36)
Substituting the complexity scalings $C_{\mathrm{BJ}}=\mathcal{O}\left(NL^{2}\right)$ and $C^{\mathrm{it}}=\mathcal{O}\left(N^{2}\right)$, we obtain
$\Delta k>\mathcal{O}\left(\frac{L^{2}}{N}\right).$ (37)
From (37), for a certain user-side antennas $N$, the break-even number of iterations saved scales approximately with $L^{2}$. Equivalently, for a fixed block size $L$, the required iteration savings decrease as $N$ increases. This indicates that, for large $N$ and moderate block sizes $L$, the BJ preconditioning cost is easily amortized, and FP-BJ-CG is computationally advantageous. Conversely, for small $N$ and large $L$, the preconditioning overhead may outweigh the benefits, and FP-CG becomes preferable. Simulation results in Sec. V will provide the simulated break-even points.
Remark 5 (Case When Assumption 2 Does Not Hold).
Assumption 2 is a strong idealization. When Assumption 2 does not hold, we define $\mathbf{M}*$ as the optimal preconditioning matrix and have the following inequality [doi:10.1137/0720040, bj_nr]:
$\kappa\left(\mathbf{M}_{*}^{1/2}{\bf A}\mathbf{M}_{*}^{1/2}\right)\leq\kappa\left(\bar{\bf A}\right)\leq d\cdot\kappa\left(\mathbf{M}_{*}^{1/2}{\bf A}\mathbf{M}_{*}^{1/2}\right).$ (38)
As shown in (38), when Assumption 2 is relaxed, the BJ preconditioning no longer yields the optimal condition number. Instead, its condition number lies within a bounded narrow interval $\left[\kappa\left(\mathbf{M}_{*}^{1/2}{\bf A}\mathbf{M}_{*}^{1/2}\right),d\cdot\kappa\left(\mathbf{M}_{*}^{1/2}{\bf A}\mathbf{M}_{*}^{1/2}\right)\right]$.